the number of diagonal matrix of order n for which A^3=A.

Grade:11

1 Answers

Sher Mohammad IIT Delhi
10 years ago
(A^3-A)=0
A(A^2-I)=0
imply A^2=I,
means only possible diagonal entries are 1 and -1, so for order n, 2*n matrix are possible.

sher mohammad
b.tech, iit delhi